# Finetuning Large Language Model as an Effective Symbolic Regressor

### v1

@@ Abstract
Deriving governing equations from observational data, known as Symbolic Regression (SR), is a cornerstone of scientific discovery. Large Language Models (LLMs) have shown promise in this task by leveraging their vast cross-disciplinary scientific knowledge. However, existing LLM-based methods primarily rely on direct inference or prompt engineering, often requiring excessive inference iterations to converge on correct formulas or failing to treating complex equation targets. These limitations in effectiveness and generalization stem from an inherent tension between pre-trained LLMs’ proficiency in approximate reasoning and the high-precision demands of SR tasks. To bridge this gap, we propose to fine-tune LLMs for enhanced SR capability. Yet, the absence of dedicated datasets for SR-oriented fine-tuning remains a critical barrier. We thus introduce SymbArena, specifically engineered to optimize LLMs for SR. This benchmark comprises 148,102 diverse equations formulated as corpora of 1.83 billion tokens for LLM utilization, enabling effective training and inference. Further, SymbArena proposes a heuristics metric to precisely quantify form-level consistency, going beyond existing SR numerical-oriented evaluation strategies. With this benchmark, we explore mainstream LLM fine-tuning techniques for SR tasks and establish SymbolicChat, a simple yet effective LLM-based SR strong baseline. Experimental results validate SymbolicChat as the first LLM to exceed traditional numerical methods in both numerical precision and symbolic form accuracy, outperforming the second-best LLM baseline with improvements of 2-fold gains in $R^{2}$ score and 8.37% in form-level consistency score.
Introduction
@@
Symbolic Regression (SR), which aims to derive the underlying governing equations from observational data, is a fundamental task in scientific discovery (Wang et al. 2023; Cornelio et al. 2023). The main challenge of the SR task is the difficulty of optimization for the regressed symbolic equation. Early solutions are usually based on numeral optimization, such as genetic programming (Schmidt and Lipson 2009; Mei et al. 2022), discovering equation heuristically through evolutionary algorithms. With the development of deep learning, task-specific supervised learning methods achieve satisfactory progress in both effectiveness and accuracy (Biggio et al. 2021; Kamienny et al. 2022; Shojaee et al. 2023). Yet, obviously, a good symbolic regressor requires complex background knowledge to support equation discovery. Hence, in the LLM period, LLM-based SR has attracted more and more attentions.
@@
Figure 1: Comparison of our method against baselines on numerical accuracy $R^{2}$ and form-level consistency. (a) and (c) show the average performance across the entire dataset, while (b) and (d) detail a representative case study. The Random baseline was generated by prompting an LLM to infer the equation without using the corpora. (a) and (c) reveal that the iterative approach of baseline methods is largely ineffective, which is essentially equivalent to exhaustive search rather than inference. Their performance curves exhibit minimal marginal gain, plateauing at a level only slightly above the Random baseline. This confirms that this spurious inference process offers no significant statistical advantage with more iterations. Conversely, our model achieves significant results at the first time inference and reason out the correct equation faster as shown in (b) and (d). In contrast, all baselines struggle to make meaningful progress through iteration. In addition, the statistically significant marginal gains shown in (a) and (c) also demonstrate the high effectiveness and generalization of our model.
@@
Existing LLM-based SR approaches, based on direct inference or prompt engineering, operate by iteratively generating a set of inferred equations and retaining only the best-fitting equation. Although the promising of the LLM having rich scentific prior, the LLMs often suffer from limitations in effectiveness and generalization, as shown in Fig. 1. The main reason for this phenomenon comes from an inherent tension between pre-trained LLMs’ proficiency in approximate reasoning and the high-precision demands of SR tasks. Prevailing general-purpose Large Language Models (LLMs) are trained to generating ambiguous but diverse outputs to fulfilled human perception. However, the ambiguous and inaccuracy in the answer induces catastrophic precision degradation in the SR, as minor symbolic deviations (e.g., operator substitution ×→+ or coefficient alteration 2.0→2.1) propagate into error accumulation, leading to physical error of derived equations. To bridge this gap, fine-tuning is a optimal strategy which could introduce task-specific constraints and accurate knowledge (Hu et al. 2022; Bai et al. 2022; Rafailov et al. 2023; Han et al. 2024). Yet, the absence of dedicated datasets for SR-oriented fine-tuning remains a barrier.
@@
To address this challenge, in this paper, we develop a new symbolic regression dataset and benchmark, termed SymbArena, to facilitate LLM-based SR fine-tuning. It comprises 148,102 diverse equations, which collectively form a massive corpus of 1.83 billion tokens. The entire dataset covers a broad spectrum of mathematical structures and complexity levels and each equation is accompanied by a task instruction, its corresponding numerical data.To operate fair evaluation, we partition the dataset according to equation skeletons, preventing potential information leakage on the equation structural level across the training and test sets. Except that, for sufficient evaluation, the SymbArena assesses symbolic regression models on two crucial perspectives. On the one hand, the SymbArena introduces numerical-level evaluation metrics, error term $R^{2}$ and tolerance-based accuracy ($\text{Acc}_{\tau}$), following existing SR benchmarks to quantify data fitting fidelity, essentially the accuracy of the dependent variable output by the regressed formula. On the other hand, the SymbArena proposes to evaluate form-level consistency, including a LLM-based metric and a well-designed heuristic metric. The former utilizes LLM to measure the form similarity between the model output and reference, while the latter are LLM-independent, measuring substring similarity between predicted and ground-truth mathematical expression strings in their coefficient-abstracted canonical forms. This novel metric compensates for the limitations of conventional numerical metrics, where erroneous formulations artificially depress fitting errors through over-optimized coefficients, thereby masking the true discrepancies.
@@
Building upon this dataset, we delve into mainstream LLM fine-tuning techniques on the SR tasks and conclude as a SymbolicChat, a strong LLM-based baseline specifically tailored for symbolic regression tasks. SymbolicChat leverages a novel Form-GRPO to operate the reinforcement fine-tuning (RFT) based on a series of manually designed form reward rules to guide structure-aware generation and improves symbolic fidelity, which goes beyond traditional model-based reward schemes on both effectiveness and generalization, as shown in Fig. 1. Experimental results validate our SymbolicChat as the first LLM to exceed traditional numerical methods in both tolerance-based accuracy and form-level consistency, Furthermore, SymbolicChat significantly outperforms the next-best LLM baseline, with 2-fold improvements of $R^{2}$ score and 8.37% in form-level consistency score.
@@
  • We propose SymbolicChat, a strong LLM-based baseline to handle symbolic regression tasks, which utilizes a novel Form-GRPO to achieve a new state of the art.
Related work
Traditional Symbolic Regression
Symbolic Regression (SR) is an emerging machine learning technique that aims to directly infer concise, interpretable mathematical equations from data to reveal underlying regularities. The field was initially dominated by genetic programming, with representative works on discovering physical conservation laws from experimental data (Schmidt and Lipson 2009). To improve efficiency and address the scalability issues of GP, methods based on sparse regression were subsequently developed , among which the representative SINDy (Brunton, Proctor, and Kutz 2016; Rudy et al. 2017), sparse identification of nonlinear dynamics, identifies dynamical equations by constructing a function library and using sparse optimization.
In recent years, deep learning has brought new breakthroughs to symbolic regression, with its applications spanning multiple levels of the equation discovery process. Researchers utilize reinforcement learning to treat equation generation as a sequential decision problem for policy optimization (Li et al. 2025b), or apply transformer architectures for large-scale pre-training, which enables the rapid generation of candidate equation skeletons from input data (Biggio et al. 2021; Kamienny et al. 2022; Shojaee et al. 2023). Furthermore, deep learning is also used to improve the foundations of equation discovery, for instance by combining autoencoders to learn low-dimensional latent representations of data (Champion et al. 2019; Li et al. 2025a), thereby discovering simple dynamical models in a more optimal coordinate system. At the same time, hybrid methods also show great potential, such as using models like RNNs to generate high-quality initial populations to improve the starting point and efficiency of traditional genetic search (Mundhenk et al. 2021; Landajuela et al. 2022).
@@
LLM for Symbolic Regression
To incorporate physical knowledge into the equation discovery process, recent symbolic regression methods based on Large Language Models (LLMs) have been proposed, which use LLMs as hypothesis generators to iteratively optimize equations via data-driven feedback (Li et al. 2025a). In terms of implementation, the work of LLM-SR (Shojaee et al. 2024) explores various iterative optimization strategies: either using the LLM as a black-box optimizer for self-improvement based on historical samples, or instructing it to perform evolutionary operations like crossover and mutation for a global search. SGA (Ma et al. 2024) proposes a bilevel optimization paradigm, with the LLM performing discrete hypothesis generation (e.g., physical equations) at the upper level, while a differentiable simulator handles the optimization of continuous parameters at the lower level. LaSR (Grayeli et al. 2024) takes a different approach by using an LLM to learn and evolve a library of abstract textual concepts, which are then used to guide and accelerate the evolutionary search for equations. However, a new benchmark (Shojaee et al. 2025) points out that when models are strictly prevented from reciting known equations, the performance of general-purpose LLMs drops sharply. This highlights the limitations of relying solely on inference and emphasizes the importance of model adaptation with domain-specific data.
SymbArena
@@
Data generation
The entire dataset $D$ contains several equations and corresponding data points, written as $D=\{(d_{i},f_{i})\}^{N}_{i=1}$, where $f_{i}$ is the equation form. $d_{i}\in\mathcal{R}^{K\times(C_{i}+1)}$ is the data matrix containing $K$ groups of data pair and each pair consists of $C_{i}$ independent variable values and one dependent variable value. To formulate such $D$, we follow two steps: 1) Generating equation $f_{i}$ and 2). Sample data points as $d_{i}$.
Equation Generation. The entire equation generation process is shown in Fig. 2. To generate equations in a program way, we represent each equation $f_{i}$ as a tree structure (Lample and Charton 2019; Kamienny et al. 2022), where each tree node represents either a mathematical operator or a terminal symbol. Then, the equation is generated following:
@@
  • Unique and Complexity Check: For each synthesized equation, we check its uniqueness by measuring the skeleton string similarity to make sure each equation is as unique as possible. Also, we check equation complexity by scanning layer number of each tree, and filter out over-simple (lower than 4) or over-complex (larger than 8) equations, make sure data do not overload model learning but also are not too simplistic for real application.
@@
Data Statistics. As shown in Tab. 1, SymbArena is compared with several widely adopted SR benchmarks, including Nguyen, R rational, SRbench, and LLM-SRbench. SymbArena contains 148,102 equations, significantly exceeding the size of existing datasets. Unlike the other datasets, SymbArena is designed for both training and testing. Among these, 11,000 equations are used for training and 512 for testing, allowing a standardized evaluation pipeline. In addition, SymbArena supports both traditional SR methods and LLM-based approaches, providing a unified platform for evaluating diverse SR paradigms.
@@
Figure 3: Overview of the proposed SymbolicChat framework. The training process consists of three stages: (1) instruction tuning on 90% of the data, (2) rule-based reward tuning (Form-GRPO) on the remaining 10%, and (3) iterative inference using SymbolicChat to refine symbolic equations.
Metric
@@
    Numerical Accuracy: Measures how closely the predicted outputs match the ground-truth values. Following the evaluation protocols proposed in prior studies (Kamienny et al. 2022; Biggio et al. 2021; La Cava et al. 2021), we adopt two widely used metrics to assess numerical accuracy: the coefficient of determination ($R^{2}$) and tolerance-based accuracy ($\text{Acc}_{\tau}$), which reports the proportion of predictions within a fixed absolute error tolerance $\tau$, given by:
    $R^{2}=1-\frac{\sum_{i}^{N_{\text{test }}}\left(y_{i}-\hat{y}_{i}\right)^{2}}{\sum_{i}^{N_{\text{test }}}\left(y_{i}-\bar{y}\right)^{2}},$ (1)
    $\mathrm{Acc}_{\tau}=\mathds{1}\left(\max_{1\leq i\leq N_{\text{test}}}\left|\frac{\hat{y}_{i}-y_{i}}{y_{i}}\right|\leq\tau\right),$ (2)
    where $\hat{y}_{i}$ and $y_{i}$ are the dependent outputs generated by feeding the same data into the predicted and ground-truth equations.
@@
    Form-level Consistency: Evaluates whether the skeleton structure of the predicted equation is consistent with the ground truth. To achieve this, we propose a comprehensive, multi-faceted Form-level Consistency metric. Specifically, we first extract the formula skeletons from the original equations. We then quantify the similarity by first decomposing each skeleton into a vector of six key structural features: operators, functions, variables, constants, structural pattern, and complexity. The similarity for each feature pair is calculated individually using methods like the Jaccard index for sets and relative ratios for counts. The final form-level consistency score, $S_{\text{struct}}$, is then computed as a weighted sum of these individual component similarities:
    $S_{\text{struct}}=\frac{1}{6}\sum_{i}\text{Sim}(\hat{f}_{i},f_{i}),$ (3)
    The details of the similarity function can be seen in the Supplementary. where $\hat{f}_{i}$ and $f_{i}$ are the predicted and ground-truth equations, respectively.
@@
    Moreover, we leverage GPT-4o as a semantic adjudicator (Achiam et al. 2023) to serve as a scalable, cost-effective proxy for human expert evaluation, providing a holistic consistency score from 0 to 1 based on the structural correspondence between the predicted and ground-truth skeletons. By introducing these two flexible scoring mechanism, we address a key limitation of previous work (La Cava et al. 2021; Shojaee et al. 2025), which often relies on a binary notion of correctness (i.e., equivalent or not) and thus fails to capture the fine-grained form-level consistency between equations. Such graded consistency measures can offer deeper insights into where symbolic regression models succeed or fail.
Method
@@
In this section, we propose a strong LLM-based SR baseline, SymbolicChat. Its main idea is to fine-tune the SR data for the LLM to drive highly effective iterative inference. The entire pipeline is shown in Figure 3. The model is first fine-tuned by instruction tuning, following a reinforcement fine-tuning by Form-GRPO. During inference, the model utilizes iterative inference to refine outputs, leading to accurate equations.
Instruction Tuning

### v2

@@ Abstract
Deriving governing equations from observational data, known as Symbolic Regression (SR), is a cornerstone of scientific discovery. Large Language Models (LLMs) have shown promise in this task by leveraging their vast cross-disciplinary scientific knowledge. However, existing LLM-based methods primarily rely on direct inference or prompt engineering, often requiring excessive inference iterations to converge on correct formulas or failing to treating complex equation targets. These limitations in effectiveness and generalization stem from an inherent tension between pre-trained LLMs’ proficiency in approximate reasoning and the high-precision demands of SR tasks. To bridge this gap, we propose to fine-tune LLMs for enhanced SR capability. Yet, the absence of dedicated datasets for SR-oriented fine-tuning remains a critical barrier. We thus introduce SymbArena, specifically engineered to optimize LLMs for SR. This benchmark comprises over 148,000 diverse equations formulated as corpora of 1.83 billion tokens for LLM utilization, enabling effective training and inference. Further, to ensure a more comprehensive and fair evaluation, SymbArena proposes a heuristics metric to precisely quantify form-level consistency, going beyond existing SR numerical-oriented evaluation strategies. With this benchmark, we explore mainstream LLM fine-tuning techniques for SR tasks and establish Symbolic-R1, a simple yet effective LLM-based SR strong baseline. Experimental results validate Symbolic-R1 as the first LLM to exceed traditional numerical methods in both numerical precision and symbolic form accuracy, outperforming the second-best LLM baseline with improvements of 2-fold gains in $R^{2}$ score and 10.3% in form-level consistency score.
[…]
1 Introduction
@@
Symbolic Regression (SR), which aims to derive the underlying governing equations from observational data, is a fundamental task in scientific discovery (Wang et al., 2023; Cornelio et al., 2023). The main challenge of the SR task is the difficulty of optimization for the regressed symbolic equation. Early solutions are usually based on numeral optimization, such as genetic programming (Schmidt & Lipson, 2009; Mei et al., 2022), discovering equation heuristically through evolutionary algorithms. With the development of deep learning, task-specific supervised learning methods achieve satisfactory progress in both effectiveness and accuracy (Biggio et al., 2021; Kamienny et al., 2022; Shojaee et al., 2023). Yet, obviously, a good symbolic regressor requires complex background knowledge to support equation discovery. Hence, in the LLM period, LLM-based SR has attracted more and more attentions.
@@
Figure 1: Comparison of our method against baselines on numerical accuracy $R^{2}$ and form-level consistency, showing both average performance (a, c) and a representative case study (b, d). The results reveal that the iterative approach of baseline methods is largely ineffective, with performance plateauing just above a random baseline. This suggests their process is closer to an exhaustive search than true inference. In contrast, our model achieves significant results in its first inference and reasons out the correct equation much faster (b, d). The substantial gains shown in (a) and (c) further confirm our model’s high effectiveness and generalization.
@@
Current LLM-based SR approaches, primarily leveraging direct inference or prompt engineering, operate by iteratively generating a set of candidate equations and retaining only the best-fitting one. Although the promising of the LLM having rich scentific prior, the LLMs often suffer from limitations in effectiveness and generalization, as shown in Fig. 1. The main reason for this phenomenon comes from an inherent tension between pre-trained LLMs’ proficiency in approximate reasoning and the high-precision demands of SR tasks. Prevailing general-purpose Large Language Models (LLMs) are trained to generating ambiguous but diverse outputs to fulfilled human perception. However, the ambiguous and inaccuracy in the answer induces catastrophic precision degradation in the SR, as minor symbolic deviations (e.g., operator substitution ×→+ or coefficient alteration 2.0→2.1) propagate into error accumulation, leading to physical error of derived equations. To bridge this gap, fine-tuning is a optimal strategy which could introduce task-specific constraints and accurate knowledge (Hu et al., 2022; Bai et al., 2022; Rafailov et al., 2023; Han et al., 2024). Yet, the absence of dedicated datasets for SR-oriented fine-tuning remains a barrier.
@@
To address this challenge, in this paper, we develop a new symbolic regression dataset and benchmark, termed SymbArena, to facilitate LLM-based SR fine-tuning. It comprises over 148,000 diverse equations, which collectively form a massive corpus of 1.83 billion tokens. The entire dataset covers a broad spectrum of mathematical structures and complexity levels and each equation is accompanied by a task instruction, its corresponding numerical data.To ensure a fair evaluation free from pre-training data contamination, we first confirm the novelty of our synthetically generated equations and then partition the dataset according to equation skeletons, preventing potential information leakage on the equation structural level across the training and test sets. Except that, for sufficient evaluation, the SymbArena assesses symbolic regression models on two crucial perspectives. On the one hand, the SymbArena introduces numerical-level evaluation metrics, error term $R^{2}$ and tolerance-based accuracy ($\text{Acc}_{\tau}$), following existing SR benchmarks to quantify data fitting fidelity, essentially the accuracy of the dependent variable output by the regressed formula. On the other hand, the SymbArena proposes to evaluate form-level consistency, including a LLM-based metric and a well-designed heuristic metric. The former utilizes LLM to measure the form similarity between the model output and reference, while the latter are LLM-independent, measuring substring similarity between predicted and ground-truth mathematical expression strings in their coefficient-abstracted canonical forms. This novel metric compensates for the limitations of conventional numerical metrics, where erroneous formulations artificially depress fitting errors through over-optimized coefficients, thereby masking the true discrepancies.
@@
Building upon this dataset, we delve into mainstream LLM fine-tuning and inference techniques on the SR tasks and conclude as a Symbolic-R1, a strong LLM-based baseline specifically tailored for symbolic regression tasks. Symbolic-R1 leverages a novel Form-GRPO to operate the reinforcement fine-tuning (RFT) based on a series of manually designed form reward rules to guide structure-aware generation and improves symbolic fidelity. Subsequently, we introduce a Hypothesis–Experiment–Revision (HER) inference framework, an interative strategy designed to yield more reliable equations. By circumventing the limitations of traditional model-based reward schemes, Symbolic-R1 achieves superior effectiveness and generalization, as shown in Fig. 1. Experimental results validate our Symbolic-R1 as the first LLM to exceed traditional numerical methods in both tolerance-based accuracy and form-level consistency, Furthermore, Symbolic-R1 significantly outperforms the next-best LLM baseline, with 2-fold improvements of $R^{2}$ score and 10.3% in form-level consistency score.
@@
  • We propose Symbolic-R1, a strong LLM-based baseline to handle symbolic regression tasks, which utilizes a novel Form-GRPO with a Hypothe- sis–Experiment–Revision inference framework to achieve a new state of the art.
@@
2 SymbArena
@@
2.1 Data generation
The entire dataset $D$ contains several equations and corresponding data points, written as $D=\{(d_{i},f_{i})\}^{N}_{i=1}$, where $f_{i}$ is the equation form. $d_{i}\in\mathcal{R}^{K\times(C_{i}+1)}$ is the data matrix containing $K$ groups of data pair, each of which consists of $C_{i}$ independent variable values and one dependent variable value. To formulate such $D$, we follow two steps: 1) Generating equation $f_{i}$ and 2). Sample data points as $d_{i}$.
Equation Generation. The entire equation generation process is shown in Fig. 2. To generate equations in a program way, we represent each equation $f_{i}$ as a tree structure (Lample & Charton, 2019; Kamienny et al., 2022), where each tree node represents either a mathematical operator or a terminal symbol. Then, the equation is generated following:
@@ 2.1 Data generation
  • Unique and Complexity Check: For each synthesized equation, we check its uniqueness by measuring the skeleton string similarity to make sure each equation is as unique as possible. Also, we check equation complexity by scanning layer number of each tree, and filter out over-simple (lower than 4) or over-complex (larger than 12) equations, make sure data do not overload model learning but also are not too simplistic for real application.
@@
Data Statistics. As shown in Tab. 1, SymbArena is compared with several widely adopted SR benchmarks, including Nguyen (Uy et al., 2011), R rational (Krawiec & Pawlak, 2013), SRbench (La Cava et al., 2021), and LLM-SRbench (Shojaee et al., 2025). SymbArena contains 148,102 equations, significantly exceeding the size of existing datasets. Unlike the other datasets, SymbArena is designed for both training and testing. Among these, 147,590 equations are used for training and 512 for testing, allowing a standardized evaluation pipeline. In addition, SymbArena supports both traditional SR methods and LLM-based approaches, providing a unified platform for evaluating diverse SR paradigms.
@@
Figure 3: Overview of the proposed Symbolic-R1 framework. It consists of a two-stage training phase (1. Instruction Tuning; 2. Reinforcement Tuning with Form-GRPO) followed by an inference phase, where a Hypothesis–Experiment–Revision loop is used to refine equations.
2.2 Metric
@@
    Numerical Accuracy: Measures how closely the predicted outputs match the ground-truth values. Following the evaluation protocols proposed in prior studies (Kamienny et al., 2022; Biggio et al., 2021; La Cava et al., 2021), we adopt two widely used metrics to assess numerical accuracy: the coefficient of determination ($R^{2}$) and tolerance-based accuracy ($\text{Acc}_{\tau}$), which reports whether the worst-case relative error is within tolerance $\tau$, given by:
    $R^{2}=1-\frac{\sum_{i=1}^{N_{\text{test }}}\left(f(x_{i})-\hat{f}(x_{i})\right)^{2}}{\sum_{i=1}^{N_{\text{test }}}\left(f(x_{i})-\overline{f(x_{i})}\right)^{2}},$ (1)
    $\mathrm{Acc}_{\tau}=\mathds{1}\left(\max_{1\leq i\leq N_{\text{test}}}\left|\frac{\hat{f}(x_{i})-f(x_{i})}{f(x_{i})}\right|\leq\tau\right),$ (2)
    where $\hat{f}(x_{i})$ and $f(x_{i})$ are the dependent outputs generated by feeding the same data into the predicted and ground-truth equations.
@@
    Form-level Consistency: Evaluates whether the skeleton structure of the predicted equation is consistent with the ground truth. To achieve this, we propose a comprehensive, multi-faceted Form-level Consistency metric. Specifically, we first extract the formula skeletons from the original equations. We then quantify the similarity by first decomposing each skeleton into a vector of six key structural features: operators, functions, variables, constants, structural pattern, and complexity. The similarity for each feature pair is calculated individually using methods like the Jaccard index for sets and relative ratios for counts. The final form-level consistency score, $S_{\text{struct}}$, is then computed as the average of these individual component similarities:
    $S_{\text{struct}}=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\text{Sim}(\hat{f}{i},f{i}),$ (3)
@@
    The details of the similarity function can be seen in the Supplementary. where $\hat{f}_{i}$ and $f_{i}$ are the predicted and ground-truth equations, respectively. Moreover, we leverage GPT-4o as a semantic adjudicator (Achiam et al., 2023) to serve as a scalable, cost-effective proxy for human expert evaluation, providing a holistic consistency score from 0 to 1 based on the structural correspondence between the predicted and ground-truth skeletons. By introducing these two flexible scoring mechanism, we address a key limitation of previous work (La Cava et al., 2021; Shojaee et al., 2025), which often relies on a binary notion of correctness (i.e., equivalent or not) and thus fails to capture the fine-grained form-level consistency between equations. Such graded consistency measures can offer deeper insights into where symbolic regression models succeed or fail.
3 Method
@@
In this section, we propose a strong LLM-based SR baseline, Symbolic-R1. Its main idea is to fine-tune the SR data for the LLM to drive highly effective iterative inference. The entire pipeline is shown in Figure 3. The model is first fine-tuned by instruction tuning (Sec. 4.1), followed by reinforcement fine-tuning with Form-GRPO (Sec. 4.2). During inference, the model utilizes the iterative reflection framework (Sec. 4.3) to refine outputs, leading to accurate equations.
3.1 Instruction Tuning